# Alon’s transmitting problem and multicolor Beck–Spencer Lemma

### v1

@@ 1. Introduction
$b(H(n,q))=\left\lfloor\left(1-\tfrac{1}{q}\right)n+\tfrac{q+1}{2}\right\rfloor$.
@@ 2. Proofs
We want to restate Lemma 3 in the form of Lemma 1. For this, we need to assign a vector to a vertex in $H(n,q)$ so that the vector behaves nicely with respect to the inner product and the Hamming distance. To this end, we use vector-coloring introduced by Doerr and Srivastav in [5]. For a $qn$-dimensional vector ${\bm{z}}=(z_{1},\ldots,z_{qn})\in{\mathbb{R}}^{qn}$, let $\bar{z}_{j}:=(z_{q(j-1)+1},z_{q(j-1)+2},\ldots,z_{qj})\in{\mathbb{R}}^{q}$, and we also write ${\bm{z}}=(\bar{z}_{1},\ldots,\bar{z}_{n})$. (We write $\bar{z}$ to indicate the vector is $q$-dimensional.) For each color $i\in[q]$, we assign a vector $\bar{c}_{i}=-\frac{1}{q}\bar{1}+\bar{e}_{i}\in{\mathbb{R}}^{q}$, that is, $\bar{c}_{i}$ has $1-\frac{1}{q}$ at the $i$-th entry, and $-\frac{1}{q}$ at the other $q-1$ entries. For example, if $q=3$ then $\bar{c}_{1}=(\frac{2}{3},-\frac{1}{3},-\frac{1}{3})$, $\bar{c}_{2}=(-\frac{1}{3},\frac{2}{3},-\frac{1}{3})$, $\bar{c}_{3}=(-\frac{1}{3},-\frac{1}{3},\frac{2}{3})$. By definition the sum of all entries of $c_{i}$ equals 00. Let $Q$ be the set of color vectors, that is,
$Q=\{\bar{c}_{1},\bar{c}_{2},\ldots,\bar{c}_{q}\}.$
For each vertex $v=(\nu_{1},\ldots,\nu_{n})\in[q]^{n}$ of $H(n,q)$, we assign a vector ${\bm{z}}=(\bar{z}_{1},\ldots,\bar{z}_{n})\in{\mathbb{R}}^{qn}$ by
@@
  • $|\bar{a}\cdot\bar{x}-\bar{a}\cdot\bar{y}|\leq 1$. Moreover, if equality holds, then (i) $\bar{y}\in Q$ or (ii) there exists $i$ such that $\bar{a}=\bar{x}=\bar{c}_{i}$ and $i$-th entry of $\bar{y}$ is $-\frac{1}{q}$.
@@
Without loss of generality we may assume that $\bar{a}=\bar{c}_{1}$.
The first item follows from
@@
We will solve a system of equations satisfying these conditions according to the following procedure.
@@
Now we have $qn-\sum_{r=1}^{q}rf_{r}$ variables, and $(n-f_{q}-1)+(n-f_{q})=2n-2f_{q}-1$ equations. Thus the dimension of solution space is $(q-2)n-\sum_{r=1}^{q}rf_{r}+2f_{q}+1\geq 2f_{q}+1\geq 1$, and we have a non-trivial solution ${\bm{y}}\neq{\mathbf{0}}$. For every $\lambda\geq 0$, $\lambda{\bm{y}}$ also satisfies Eq($q$). We extend ${\bm{y}}\in{\mathbb{R}}^{[qn]\setminus F_{q}({\bm{x}}_{s})}$ to ${\bm{y}}\in{\mathbb{R}}^{qn}$ by inserting 00 at the positions in $F_{q}({\bm{x}}_{s})$. By continuously increasing $\lambda$ from 00, we get $\lambda{\bm{y}}$ such that one of the entries becomes $-\frac{1}{q}$ or $1-\frac{1}{q}$ for the first time. Using this $\lambda>0$, let ${\bm{x}}:={\bm{x}}_{s}+\lambda{\bm{y}}$. Then ${\bm{x}}$ satisfies (C1) for $1\leq i\leq n-s$, (C2) for $1\leq j\leq n$, and (C3). Update ${\bm{x}}_{s}:={\bm{x}}$, and return to the beginning of Step(s).
@@
Thus $w\not\in\Gamma_{m-1}(v_{1})\cup\cdots\cup\Gamma_{0}(v_{m})$, which means that $b(H,q)\geq m+1$. ∎
@@ 3. Concluding remarks
For the proof of Theorem 1, Alon observed the following. In Lemma 1 we have $|{\bm{a}}_{i}\cdot{\bm{x}}|=|{\bm{a}}_{i}\cdot(-{\bm{x}})|$, and so the inequality on $d(v_{i},w)$ in Lemma 2 also holds if we replace $v_{i}$ with its antipodal vertex. It would be nice if a similar extension was possible for Lemma 3, which would improve Theorem 2. For simplicity, let $q=3$ and $n=3k+1$. For $v=(x_{1},\ldots,x_{n})\in\{0,1,2\}^{n}$, let $v^{\prime}=(x_{1}+1,\ldots,x_{n}+1)$ and $v^{\prime\prime}=(x_{1}+2,\ldots,x_{n}+2)$, where addition is done in modulo 3. For vertices $u$ and $v$, let $f(u,v):=|\frac{2}{3}(3k+1)-d(u,v)|$, and $g(u,v):=\max\{f(u,v),f(u,v^{\prime}),f(u,v^{\prime\prime})\}$.
@@
For $11\leq i\leq 3k+$ 1, let $u_{i}$ be a given vertex of $H(3k+1,3)$. Then is it true that there exists a vertex $w$ such that $g(u_{i},w)<i$ for all $11\leq i\leq 3k+$ 1?
If this is true, then it implies that $b(H(3k+1,3))=2k+2$. To see this, assume that the answer to the problem is affirmative, and let vertices $v_{1},v_{2},\ldots,v_{2k+1}$ be given in $H(3k+1,3)$. As in [1], we apply the existence of $w$ to $u_{1}:=v_{2},u_{2}:=v_{3},\ldots,u_{2k}:=v_{2k+1}$. Then we have $g(v_{i},w)<i-1$ for $i=2,3,\ldots,2k+1$. Thus all of $w,w^{\prime}$ and $w^{\prime\prime}$ have distance at least $2k+2-i$ from $v_{i}$ for all $12\leq i\leq 2k+$ 1. Finally we use the fact that one of $w,w^{\prime}$, and $w^{\prime\prime}$ has distance at least $12k+$ 1 from $v_{1}$ because $d(v_{1},w)+d(v_{1},w^{\prime})+d(v_{1},w^{\prime\prime})=2(3k+1)$. Consequently, there is $w^{*}\in\{w,w^{\prime},w^{\prime\prime}\}$ such that $d(v_{i},w^{*})\geq 2k+2-i$ for all $11\leq i\leq 2k+$ 1, that is, $b(H(3k+1,3))\geq 2k+2$.
@@ Acknowledgments
The author thanks Naoki Matsumoto for stimulating discussions. The author was supported by JSPS KAKENHI Grant Number JP23K03201.

### v2

@@ Abstract.
Key words and phrases:
Hamming graph, burning number, floating variable, discrepancy
1991 Mathematics Subject Classification:
94A05, 05C35, 90C10
@@ 1. Introduction
For fixed $q\geq 3$ and large enough $n$, we have $b(H(n,q))=\left\lfloor\left(1-\tfrac{1}{q}\right)n+\tfrac{q+1}{2}\right\rfloor$.
Since $H(n,q)$ has diameter $n$, we have a trivial bound $b(H(n,q))\leq n+1$. Thus if the conjecture holds, then we need $n\geq\lfloor\frac{q(q-1)}{2}\rfloor$. On the other hand, it follows from the lower bound in Theorem 2 that $b(H(n,q))=n+1$ if $q\geq n\geq 3$, see the last section.
@@ 2. Proofs
We want to restate Lemma 3 in the form of Lemma 1. For this, we need to assign a vector to a vertex in $H(n,q)$ so that the vector behaves nicely with respect to the inner product and the Hamming distance. To this end, we use vector-coloring introduced by Doerr and Srivastav in [5]. For a $qn$-dimensional vector ${\bm{z}}=(z_{1},\ldots,z_{qn})\in{\mathbb{R}}^{qn}$, let $\bar{z}_{j}:=(z_{q(j-1)+1},z_{q(j-1)+2},\ldots,z_{qj})\in{\mathbb{R}}^{q}$, and we also write ${\bm{z}}=(\bar{z}_{1},\ldots,\bar{z}_{n})$. (We write $\bar{z}$ to indicate the vector is $q$-dimensional.) For each color $i\in[q]$, we assign a vector $\bar{c}_{i}=-\frac{1}{q}\bar{1}+\bar{e}_{i}\in{\mathbb{R}}^{q}$, that is, $\bar{c}_{i}$ has $1-\frac{1}{q}$ at the $i$-th entry, and $-\frac{1}{q}$ at the other $q-1$ entries. For example, if $q=3$ then $\bar{c}_{1}=(\frac{2}{3},-\frac{1}{3},-\frac{1}{3})$, $\bar{c}_{2}=(-\frac{1}{3},\frac{2}{3},-\frac{1}{3})$, $\bar{c}_{3}=(-\frac{1}{3},-\frac{1}{3},\frac{2}{3})$. By definition the sum of all entries of $c_{i}$ equals 00. We also have
Let $Q$ be the set of color vectors, that is, $Q=\{\bar{c}_{1},\bar{c}_{2},\ldots,\bar{c}_{q}\}$. For each vertex $v=(\nu_{1},\ldots,\nu_{n})\in[q]^{n}$ of $H(n,q)$, we assign a vector ${\bm{z}}=(\bar{z}_{1},\ldots,\bar{z}_{n})\in{\mathbb{R}}^{qn}$ by
@@
  • $|\bar{a}\cdot\bar{x}-\bar{a}\cdot\bar{y}|\leq 1$. Moreover, if equality holds, then (i) $\bar{y}\in Q$ or (ii) there exists $i$ such that $\bar{a}=\bar{x}=\bar{c}_{i}$ and the $i$-th entry of $\bar{y}$ is $-\frac{1}{q}$.
@@
Without loss of generality we may assume that $\bar{a}=\bar{c}_{1}$. The first item follows from (1).
@@
We will repeatedly solve a system of equations satisfying these conditions according to the following procedure.
@@
Now we have $qn-\sum_{r=1}^{q}rf_{r}$ variables, and $(n-f_{q}-1)+(n-f_{q})=2n-2f_{q}-1$ equations. Thus the dimension of solution space is $(q-2)n-\sum_{r=1}^{q}rf_{r}+2f_{q}+1\geq 2f_{q}+1\geq 1$, and we have a non-trivial solution ${\bm{y}}\neq{\mathbf{0}}$. For every $\lambda\geq 0$, $\lambda{\bm{y}}$ also satisfies Eq($q$). We extend ${\bm{y}}\in{\mathbb{R}}^{qn-\sum_{r=1}^{q}rf_{r}}$ to ${\bm{y}}\in{\mathbb{R}}^{qn}$ by inserting 00 at the positions of the fixed variables. By continuously increasing $\lambda$ from 00, we get $\lambda{\bm{y}}$ such that one of the entries becomes $-\frac{1}{q}$ or $1-\frac{1}{q}$ for the first time. Using this $\lambda>0$, let ${\bm{x}}:={\bm{x}}_{s}+\lambda{\bm{y}}$. Then ${\bm{x}}$ satisfies (C1) for $1\leq i\leq n-s$, (C2) for $1\leq j\leq n$, and (C3). Update ${\bm{x}}_{s}:={\bm{x}}$, and return to the beginning of Step(s).
@@
Thus $w\not\in\Gamma_{m-1}(v_{1})\cup\cdots\cup\Gamma_{0}(v_{m})$, which means that $b(H(n,q))\geq m+1$. ∎
@@ 3. Concluding remarks
In this subsection, we also write $b(n,q)$ to mean $b(H(n,q))$.
3.1. The case $q\geq n$
Recall that $H(n,q)$ has diameter $n$, so $b(n,q)\leq n+1$ for all $n,q$.
Claim 3.
Let $q\geq 3$ and $s\leq n$. If $b(n,q-1)\geq s$ then $b(n,q)\geq s+1$.
Proof.
Suppose that $b(n,q-1)\geq s$ but $b(n,q)\leq s$. Then we have a burning sequence for $H(n,q)$ of length $s$, starting from, without loss of generality, $v_{1}=q{\bm{1}}$. Since $[q]^{n}\setminus\Gamma_{s-1}(v_{1})\supset[q-1]^{n}$, we need to cover $H(n,q-1)$ in less than $s$ rounds, which is impossible because $b(n,q-1)\geq s$. ∎
Proposition.
If $q\geq n\geq 3$ then $b(H(n,q))=n+1$.
Proof.
Since $b(n,q+1)\geq b(n,q)$, it suffices to show that $b(n,n)=n+1$. By Theorem 2 we have
(5) $b(n,n-2)\geq\lfloor\left(1-\tfrac{1}{n-2}\right)n\rfloor+1=\lfloor n-1-\tfrac{2}{n-2}\rfloor+1=n-1.$
Applying Claim 3 to (5) twice, we get $b(n,n)\geq n+1$. ∎
3.2. Possible extension of Lemma 3
For the proof of Theorem 1, Alon observed the following. In Lemma 1 we have $|{\bm{a}}_{i}\cdot{\bm{x}}|=|{\bm{a}}_{i}\cdot(-{\bm{x}})|$, and so the inequality on $d(v_{i},w)$ in Lemma 2 also holds if we replace $v_{i}$ with its antipodal vertex. It would be nice if a similar extension was possible for Lemma 3, which would improve Theorem 2. For simplicity, let $q=3$ and $n=3k+1\geq 4$ be fixed. For $v=(x_{1},\ldots,x_{n})\in\{0,1,2\}^{n}$, let $v^{\prime}=(x_{1}+1,\ldots,x_{n}+1)$ and $v^{\prime\prime}=(x_{1}+2,\ldots,x_{n}+2)$, where addition is done in modulo 3. For vertices $u$ and $v$, let $f(u,v):=|\frac{2}{3}(3k+1)-d(u,v)|$, and $g(u,v):=\max\{f(u,v),f(u,v^{\prime}),f(u,v^{\prime\prime})\}$.
@@ 3.2. Possible extension of Lemma 3
For $11\leq i\leq 3k+$ 1, let $u_{i}$ be a given vertex of $H(3k+1,3)$ on $\{0,1,2\}^{n}$. Then is it true that there exists a vertex $w$ such that $g(u_{i},w)<i$ for all $11\leq i\leq 3k+$ 1?
If this is true, then it implies that $b(3k+1,3)=2k+2$. To see this, assume that the answer to the problem is affirmative, and let vertices $v_{1},v_{2},\ldots,v_{2k+1}$ be given in $H(3k+1,3)$. As in [1], we apply the existence of $w$ to $u_{1}:=v_{2},u_{2}:=v_{3},\ldots,u_{2k}:=v_{2k+1}$. Then we have $g(v_{i},w)<i-1$ for $i=2,3,\ldots,2k+1$. Thus all of $w,w^{\prime}$ and $w^{\prime\prime}$ have distance at least $2k+2-i$ from $v_{i}$ for all $12\leq i\leq 2k+$ 1. Finally we use the fact that one of $w,w^{\prime}$, and $w^{\prime\prime}$ has distance at least $12k+$ 1 from $v_{1}$ because $d(v_{1},w)+d(v_{1},w^{\prime})+d(v_{1},w^{\prime\prime})=2(3k+1)$. Consequently, there is $w^{*}\in\{w,w^{\prime},w^{\prime\prime}\}$ such that $d(v_{i},w^{*})\geq 2k+2-i$ for all $11\leq i\leq 2k+$ 1, that is, $b(3k+1,3)\geq 2k+2$.
@@ Acknowledgments
The author thanks Noga Alon for valuable comments, in particular, for pointing out an error in the conjecture in the earlier version, and Naoki Matsumoto for stimulating discussions. The author was supported by JSPS KAKENHI Grant Number JP23K03201.